OpenStudy (baby456):

Write the equation of the line in slop -intercept form given the following inforamatation. the line passes through the point P and Q. 15. y-9=1/2(x+6) 16. P(11,8) Q(-1,8)

OpenStudy (baby456):

13. !4x-2y+9=0 need help with that also

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ P&({\color{red}{ 11}}\quad ,&{\color{blue}{ 8}})\quad Q&({\color{red}{ -1}}\quad ,&{\color{blue}{ 8}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

y - 9 = 1/2(x + 6) -- distribute through the parenthesis y - 9 = 1/2x + 3 -- add 9 to both sides y = 1/2x + 3 + 9 -- combine like terms y = 1/2x + 12 ==> slope intercept form ========================= P = (11,8) x1 = 11 and y1 = 8 Q = (-1,8) x2 = -1 and y2 = 8 slope(m) = (y2 - y1) / (x2 - x1) now we sub and find the slope slope(m) = (8-8) / (-1-11) slope(m) = 0/-12 slope = 0 (horizontal line) y = 0x + 8 y = 8 ========================== 14x - 2y + 9 = 0 14x -2y = -9 -2y = -14x - 9 -- divide by -2 y = -14/-2x - 9/-2 y = 7x + 9/2 ==> slope intercept form any questions ?

OpenStudy (baby456):

got it thanka

OpenStudy (anonymous):

good to hear :)